OpenStudy (anonymous):

Simplify. 17 – 6 × 10 ¸ 2 + 12

OpenStudy (anonymous):

is this one problem?

OpenStudy (ghazi):

2+12=14

OpenStudy (ghazi):

17-6*12 multiplication first so it's 17-72= -55

OpenStudy (shivaniits):

but i think it is 17-6*10 so one could go for 17-60=43...!!